# Course Allocation with Credits via Stable Matching

### v1

@@ 5.2 With lower quotas
Until now we have assumed that courses do not have a minimum number of students that they require in order to run, which are represented by lower quotas. However, this is a very natural requirement, present in many universities, and so we now define an extension to the Course Allocation problem involving lower quotas. A course $c_{j}$ is said to have a lower quota, denoted by $q^{-}(c_{j})$, if in a matching $M$ the number of students allocated to $c_{j}$ must be at least the lower quota, i.e., $|S_{M}(c_{j})|\geq q^{-}(c_{j})$. When courses have lower quotas, there are two variations to consider. In the first one, every course in the Course Allocation instance must have a number of students at least as large as its lower quota. The goal is to find a stable matching (where every course satisfies its lower and upper quota) for any given definition of stability. This problem is called ca-x-lq-nc-find, where x denotes the definition of stability. In the case of ca-ml, the problem is called ca-x-ml-lq-nc-find.
The second variation models the scenario where in a given matching $M$ if course $c_{j}$ does not have at least as many students as its lower quota then it may be closed, and no students are matched to it. In this case an empty matching, where every course is closed, is trivially stable, so our goal is to find a maximum size stable matching. This problem is called ca-x-lq-cl-max, where x denotes the definition of stability. In the case of ca-ml, the problem is called ca-x-ml-lq-cl-max. In this variation we do not allow for the existence of a coalition of several students and a closed course that would demand for this course to stay open (as in biro_college_2010 ), because in real-life settings lower quotas are usually too high for students to coordinate in such a fashion.
From Theorems 7 and 8 and Corollary 1 we immediately get these two results.
@@
Given an instance $I$ of ca-ml with lower quotas, ca-ps-ml-lq-nc-find, ca-c-ml-lq-nc-find and ca-fc-ml-lq-nc-find are NP-hard, even if every student takes up to $2$ credits, every course has $1$ or $2$ credits, the length of the students’ preference lists is bounded, and there are no downward-feasible constraints.
@@
ca-ps-ml-lq-cl-max, ca-c-ml-lq-cl-max and ca-fc-ml-lq-cl-max are NP-hard, even if every student takes up to $2$ credits, every course has $1$ or $2$ credits, the length of the students’ preference lists is bounded, and there are no downward-feasible constraints.
@@
Under pair-size stability, coalition stability or first-coalition stability, finding a maximum size stable matching is hard without lower quotas, so it remains hard with lower quotas. ∎
Similarly, given an instance of ca with lower quotas, determining the existence of a pair-stable matching is NP-complete, which is an immediate consequence of Theorem 3. However, in the case of master lists, we know that every instance of ca-ml has a unique pair-stable matching. Thus, in order to determine whether an instance of ca-ml with lower quotas has a pair-stable matching that respects lower quotas, we just need to run Serial Dictatorship and check whether in the resulting matching every course has enough students to fill its lower quota. Thus, we obtain the following Observation.
@@
It is NP-complete to determine if an instance of ca-ml with lower quotas and closures has a student-complete pair-stable matching, even if both students and courses are ordered according to master lists, every student takes up to $2$ credits, every course is of size $1$ or $2$ and with lower quota $1$ or $2$, and there are no downward-feasible constraints.
@@
Verifying whether a matching is student-complete and pair-stable, and respects lower quotas, can be done in polynomial time, so the problem belongs to NP. We prove NP-completeness by reducing from Exact-MM, which is NP-complete, even if $G$ is bipartite and all vertices in one of the sides have degree $2$. Let $G=(U,W,E)$ be a graph representing an instance of Exact-MM, where $U$ and $W$ are the two sets of vertices, $|U|=n_{1}$ and $|W|=n_{2}$. Every vertex in $W$ has degree $2$. Our goal is to find a maximal matching of size $K$. We build an instance $I$ of ca-ml-lq-cl, with students $S$ and courses $C$, as follows.
@@
For every $w_{i}\in W$, we create the following students: $s_{i}^{1}$, $s_{i}^{2}$, $s_{i}^{3}$, and $s_{i}^{4}$, with $T(s_{i}^{1})=T(s_{i}^{2})=T(s_{i}^{3})=T(s_{i}^{4})=2$. We also create courses $d_{i}^{1}$ and $d_{i}^{2}$, with $O(d_{i}^{1})=O(d_{i}^{2})=2$, as well as courses $b_{i}^{1}$ and $b_{i}^{2}$, with $O(b_{i}^{1})=O(b_{i}^{2})=1$. Students $s_{i}^{1}$, $s_{i}^{2}$, $s_{i}^{3}$ and $s_{i}^{4}$ belong to subset $S_{i}\subseteq S$, and courses $d_{i}^{1}$, $d_{i}^{2}$ and $d_{i}^{3}$ belong to subset $D_{i}\subseteq C$. For every $w_{i}\in W$, $s_{i}^{1}\in S^{1}\subseteq S$, $s_{i}^{2}\in S^{2}\subseteq S$, $s_{i}^{3}\in S^{3}\subseteq S$, and $s_{i}^{4}\in S^{4}\subseteq S$. Similarly, $d_{i}^{1}\in D^{1}\subseteq C$, $d_{i}^{2}\in D^{2}\subseteq C$ and $b_{i}^{1},b_{i}^{2}\in B\subseteq C$.
@@
For every $u_{j}\in U$, we create a course $c_{j}$, with $O(c_{j})=1$. Additionally, we create two courses $e$ and $f$, with $O(e)=O(f)=2$. The upper quota of every course is $1$, except for $e$, which has upper quota $n_{2}-K$, and every $b_{i}^{l}\in B$, $d_{i}^{1}\in D^{1}$ and $f$, which have upper quota $2$. The lower quota of every course is $1$, except for every $b_{i}^{l}\in B$, every $d_{i}^{1}\in D^{1}$ and $f$, which have lower quota $2$. For every $u_{j}\in U$, $c_{j}\in C^{\prime}\subseteq C$.
We also create $K$ students labelled $p_{1}$ to $p_{K}$ such that, for every $p_{k}$, $T(p_{k})=1$, and two students $q$ and $r$ such that $T(q)=T(r)=2$. For every $p_{k}$, $1\leq k\leq K$, $p_{k}\in P\subseteq S$.
The master list of students (MLS) is
$s_{1}^{1}\succ_{MLS}s_{1}^{4}\succ_{MLS}s_{1}^{2}\succ_{MLS}s_{2}^{1}\succ_{MLS}s_{2}^{4}\succ_{MLS}s_{2}^{2}\succ_{MLS}...\succ_{MLS}s_{n_{2}}^{1}\succ_{MLS}\\ \succ_{MLS}s_{n_{2}}^{4}\succ_{MLS}s_{n_{2}}^{2}\succ_{MLS}s_{1}^{3}\succ_{MLS}s_{2}^{3}\succ_{MLS}...\succ_{MLS}s_{n_{2}}^{3}\succ_{MLS}\\ \succ_{MLS}p_{1}\succ_{MLS}...\succ_{MLS}p_{K}\succ_{MLS}q\succ_{MLS}r$ (4)
@@
That is, the first students belong to subsets $S^{1}$, $S^{4}$ and $S^{2}$, such that every $s_{i}^{1}$ is followed by $s_{i}^{4}$ and then by $s_{i}^{2}$. Then we place students from $S^{3}$, and after them students from $P$, in all cases in increasing subscript order. Finally, we place students $q$ and $r$. The master list of courses (MLC) is
$b_{1}^{1}\succ_{MLC}b_{1}^{2}\succ_{MLC}b_{2}^{1}\succ_{MLC}b_{2}^{2}\succ_{MLC}...\succ_{MLC}b_{n_{2}}^{1}\succ_{MLC}b_{n_{2}}^{2}\succ_{MLC}\\ \succ_{MLC}d_{1}^{1}\succ_{MLC}d_{1}^{2}\succ_{MLC}d_{2}^{1}\succ_{MLC}d_{2}^{2}\succ_{MLC}...\succ_{MLC}d_{n_{2}}^{1}\succ_{MLC}d_{n_{2}}^{2}\succ_{MLC}\\ \succ_{MLC}c_{1}\succ_{MLC}c_{2}\succ_{MLC}...\succ_{MLC}c_{n_{1}}\succ_{MLC}e\succ_{MLC}f$ (5)
@@
That is, the first courses are those from $B$, then the next courses are from $D^{1}$ and $D^{2}$, such that for each $w_{i}\in W$, $d_{i}^{1}\succ_{MLC}d_{i}^{2}$, after them those which belong to $C^{\prime}$, and finally we have courses $e$ and $F$, in all cases in increasing subscript order.
@@
If in the Exact-MM instance $w_{i}$ is matched with $u_{j}$, then $s_{i}^{1}$ is to be matched with $c_{j}$ and $b_{i}^{1}$, $s_{i}^{2}$ with $d_{i}^{1}$, $s_{i}^{3}$ with $d_{i}^{2}$, and $s_{i}^{4}$ with $d_{i}^{1}$. If, however, $w_{i}$ is unmatched, $s_{i}^{1}$ is to be matched with $d_{i}^{1}$, $s_{i}^{2}$ with $d_{i}^{2}$, $s_{i}^{3}$ with $e$, and $s_{i}^{4}$ with $d_{i}^{1}$. Student $s_{i}^{3}$ contains courses $c_{j}$ and $c_{k}$ in her preference list to make sure that a course-complete pair-stable matching in $I$ corresponds to a maximal matching in $G$, so that if $s_{i}^{3}$ is matched with $e$, some students from $S^{1}$ or $S^{2}$ are matched with $c_{j}$ and $c_{k}$. The preference lists of agents in $P$ includes every course in $B$.
@@
Agents from $P$ are matched to the courses in $B$ that have some $s_{i}^{l}$ matched to them, which correspond to vertices $w_{i}\in G$ that are matched in $G$. The preference lists of $q$ and $r$ are:
@@
The preference list of every $c_{j}\in C^{\prime}$ depends on the vertices adjacent to $u_{j}$. For every $w_{i}$ adjacent to $u_{j}$, $c_{j}$ includes in its preference list either $s_{i}^{1}$ or $s_{i}^{2}$, as well as $s_{i}^{3}$. The order of all students is derived from the master list. If, for example, $u_{j}$ is adjacent to $w_{3}$, $w_{4}$ and $w_{5}$, the preference list of $c_{j}$ is given below. Here the superscript $l_{i}$ indicates a number that can be either $1$ or $2$. If $u_{j}$ is matched in $G$ then $c_{j}$ is matched with some agent from $S^{1}$ or $S^{2}$, otherwise it is unmatched.
@@
Figure 8: Example of an Exact-MM instance, with $K=2$, and some of the preference lists in the transformed instance of ca-ml-lq-cl.
We now show that a maximal matching $M$ of size $K$ in $G$ implies a student-complete pair-stable matching $M^{\prime}$ in $I$ that respects lower quotas. For every pair $(u_{j},w_{i})\in M$, we add to $M^{\prime}$ either the pairs $(s_{i}^{1},c_{j})$ and $(s_{i}^{1},b_{i}^{1})$, or the pairs $(s_{i}^{2},c_{j})$ and $(s_{i}^{2},b_{i}^{2})$, depending on whether $s_{i}^{1}$ or $s_{i}^{2}$ has $c_{j}$ in her preference list. If $(s_{i}^{1},c_{j})\in M^{\prime}$, then we add $(s_{i}^{2},d_{i}^{1})$ to $M^{\prime}$, otherwise we add $(s_{i}^{1},d_{i}^{1})$. We also add $(s_{i}^{3},d_{i}^{2})$ and $(s_{i}^{4},d_{i}^{1})$ to $M^{\prime}$. If $w_{i}$ has no partner in $M$, we add the following pairs to $M^{\prime}$: $(s_{i}^{1},d_{i}^{1})$, $(s_{i}^{2},d_{i}^{2})$, $(s_{i}^{3},e)$, and $(s_{i}^{4},d_{i}^{1})$. Additionally, let $j_{k}$ ($1\leq k\leq K$) be a sequence such that $w_{j_{k}}$ is matched in $M$. If $(s_{i}^{1},c_{j_{k}})\in M^{\prime}$ for some $s_{i}^{1}$, add $(p_{j_{k}},b_{i}^{1})$ to $M^{\prime}$, otherwise add $(p_{j_{k}},b_{i}^{2})$ to $M^{\prime}$. Finally, add pairs $(q,f)$ and $(r,f)$ to $M^{\prime}$. We close every course in $B$ that has no students matched to it. The matching $M^{\prime}$ is student-complete, because for every student $s\in S$, $T(s)=O(C_{M^{\prime}}(s))$. It also respects the lower quotas of every course that appears in $M^{\prime}$. Course $f$ has two students, and every open course in $B$ and $D^{1}$ has two students too. We show below that it is pair-stable.
@@
$M^{\prime}$ is a pair-stable matching.
@@
We prove the pair-size stability of $M^{\prime}$ by showing that no student can belong to a blocking pair. Every $s_{i}^{1}$ is matched with either some $c_{j}$ and $b_{i}^{1}$, which is her most preferred course, or with $d_{i}^{1}$. In the second case, $b_{i}^{1}$ is closed, so she cannot form a blocking pair with it. Similarly, every $s_{i}^{2}$ is matched with either some $c_{k}$ and $b_{i}^{2}$, her most preferred course, with $d_{i}^{1}$, or with $d_{i}^{2}$. If $s_{i}^{2}$ is matched with $b_{i}^{2}$ and $c_{k}$, or with $d_{i}^{1}$, the same reasoning applies, and if she is matched with $d_{i}^{2}$, then $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$ and $(s_{i}^{4},d_{i}^{1})\in M^{\prime}$, so there is no blocking pair containing $s_{i}^{2}$. The student $s_{i}^{4}$ is always matched with the only course in her preference list, $d_{i}^{1}$.
@@
We now consider $s_{i}^{3}$. This student is matched with either $d_{i}^{2}$, her most preferred course, or with $e$. In the second case, $(s_{i}^{2},d_{i}^{2})\in M$ and, because $M$ is maximal in $G$, $u_{j}$ and $u_{k}$ are matched with some $w_{l}$ and $w_{m}$. Thus, there exist $s_{l}^{1}$ (or $s_{l}^{2}$) and $s_{m}^{1}$ (or $s_{m}^{2}$) such that $c_{j}$ and $c_{k}$ are matched with them. Hence, $s_{i}^{3}$ cannot form a blocking pair with any other course. Moving on to students in $P$, every $p_{k}\in P$ is matched with some $b_{i}^{l}\in B$, where $l=1$ or $l=2$. Let $b_{j}^{m}$ be an open course that $p_{k}$ prefers over $b_{i}^{l}$. Then $b_{j}^{m}$ is matched with both some $s_{i}^{1}$ or $s_{i}^{2}$ and with a $p_{k}^{\prime}$ such that $p_{k}^{\prime}\succ_{MLS}p_{k}$. Thus, $p_{k}$ cannot form a blocking pair with $b_{i}^{l}$. Finally, $r$ is matched with the only course in her preference list, $f$, and $q$ is matched with $f$, too. Even though $q$ prefers $e$ over $f$, $e$ has filled its capacity, and every student enrolled in $e$ belongs to $S^{3}$, all of which appear in the master list before $q$. Therefore, no student belongs to a blocking pair. ∎
@@
Now, assume that there exists a student-complete pair-stable matching $M^{\prime}$ in $I$. We now show how to get a maximal matching $M$ of size $K$ in $G$. First, as $M^{\prime}$ is student-complete, $r$ is matched with $f$. The minimum quota of $f$ is $2$, so we have that $(q,f)\in M^{\prime}$ too. However, $q$ prefers $e$ over $f$. Thus, because $M^{\prime}$ is pair-stable, $e$ must have $n_{2}-K$ students assigned to it that are better than $q$. Hence, there are $n_{2}-K$ students from $S^{3}$ matched to $e$ in $M^{\prime}$. In order to preserve pair stability, for each $s_{i}^{3}$ matched to $e$, $M^{\prime}$ contains the pairs $(s_{i}^{1},d_{i}^{1})$ and $(s_{i}^{2},d_{i}^{2})$, as well as the pair $(s_{i}^{4},d_{i}^{1})$.

### v2

@@ 5.2 With lower quotas
A very natural requirement, present in many universities, is that courses have minimum numbers of students that they require in order to run, which are represented by lower quotas. We have not considered this restriction until now, but in this section we now define an extension to ca involving lower quotas.
A course $c_{j}$ is said to have a lower quota, denoted by $q^{-}(c_{j})$, if in a matching $M$ the number of students allocated to $c_{j}$ must be at least the lower quota, i.e., $|S_{M}(c_{j})|\geq q^{-}(c_{j})$. When courses have lower quotas, there are two variations to consider. In the first one, every course in the ca instance must have a number of students at least as large as its lower quota. The goal is to find a stable matching, where every course is assigned a number of students between its lower and upper quotas, for any given definition of stability. This problem is called ca-x-lq-nc-find, where x denotes the definition of stability. In the case of ca-ml, the problem is called ca-x-ml-lq-nc-find.
The second variation models the scenario where in a given matching $M$, if course $c_{j}$ does not have at least as many students as its lower quota, then it may be closed, and no students are matched to it in this case. In this setting an empty matching, where every course is closed, is trivially stable, so our goal is to find a maximum size stable matching. This problem is called ca-x-lq-cl-max, where x denotes the definition of stability. In the case of ca-ml, the problem is called ca-x-ml-lq-cl-max. In this variation we do not allow for the existence of a coalition of several students and a closed course where the students have an incentive to demand that this course be opened (as in Biró et al. (2010)), because in real-life settings lower quotas are usually too high for students to coordinate in such a fashion.
From Theorems 7 and 8 and Corollary 1 we immediately obtain the following two results for ca-ml with lower quotas both without and with course closures.
@@
Given an instance $I$ of ca-ml with lower quotas, each of ca-ps-ml-lq-nc-find, ca-c-ml-lq-nc-find and ca-fc-ml-lq-nc-find is NP-hard, even if every student takes up to two credits, every course has one or two credits, the length of the students’ preference lists is bounded, and there are no downward-feasible constraints.
@@
ca-ps-ml-lq-cl-max, ca-c-ml-lq-cl-max and ca-fc-ml-lq-cl-max is NP-hard, even if every student takes up to two credits, every course has one or two credits, the length of the students’ preference lists is bounded, and there are no downward-feasible constraints.
@@
Under pair-size stability, coalition stability or first-coalition stability, finding a maximum size stable matching is NP-hard without lower quotas (see Theorem 8, Corollary 1 and Theorem 7 respectively), so the problem remains hard with lower quotas. ∎
Similarly, the following result is an immediate consequence of Theorem 3:
@@
Given an instance of ca with lower quotas and no course closures, determining the existence of a pair-stable matching is NP-complete. That is, ca-p-lq-nc-find is NP-hard.
However, in the case of master lists, we know that every instance of ca-ml has a unique pair-stable matching. Thus, in order to determine whether an instance of ca-ml with lower quotas has a pair-stable matching that respects lower quotas, we just need to run Algorithm 2 and check whether in the resulting matching every course has enough students to fill its lower quota. Thus, we obtain the following Observation.
Observation 7.
@@
It is NP-complete to determine if an instance of ca-ml with lower quotas and closures has a student-complete pair-stable matching, even if both students and courses are ordered according to master lists, every student takes up to two credits, every course has one or two credits and lower quota $1$ or $2$, and there are no downward-feasible constraints. Hence, it follows that ca-p-ml-lq-cl-max is NP-hard.
@@
Verifying whether a matching is student-complete and pair-stable, and respects lower quotas, can be done in polynomial time, so the decision problem belongs to NP. We prove NP-completeness by reducing from exact-mm, which is NP-complete, even if $G$ is bipartite and all vertices in one of the sides have degree $2$ Manlove et al. (2002)(Manlove, 2013, page 133). Let $G=(U,W,E)$ be a graph and $K$ be an integer representing an instance of exact-mm, where $U$ and $W$ are the two sets of vertices in the bipartition, $|U|=n_{1}$ and $|W|=n_{2}$, and $K$ is the target size of a maximal matching. Without loss of generality assume that every vertex in $W$ has degree $2$. We build an instance $I$ of ca-ml-lq-cl, with students $S$ and courses $C$, as follows.
@@
For every $w_{i}\in W$, we create the following students: $s_{i}^{1}$, $s_{i}^{2}$, $s_{i}^{3}$, and $s_{i}^{4}$, with $T(s_{i}^{1})=T(s_{i}^{2})=T(s_{i}^{3})=T(s_{i}^{4})=2$. We also create courses $d_{i}^{1}$ and $d_{i}^{2}$, with $O(d_{i}^{1})=O(d_{i}^{2})=2$, as well as courses $b_{i}^{1}$ and $b_{i}^{2}$, with $O(b_{i}^{1})=O(b_{i}^{2})=1$. Create the following subsets: $S_{i}=\{s_{i}^{1},s_{i}^{2},s_{i}^{3},s_{i}^{4}\}$ ($1\leq i\leq n_{2}$), $D_{i}=\{d_{i}^{1},d_{i}^{2}\}$ ($1\leq i\leq n_{2}$), $S^{r}=\{s_{i}^{r}:1\leq i\leq n_{2}\}$ ($1\leq r\leq 4$), $D^{r}=\{d_{i}^{r}:1\leq i\leq n_{2}\}$ ($1\leq r\leq 2$) and $B=\{b_{i}^{r}:1\leq i\leq n_{2}\wedge 1\leq r\leq 2\}$.
@@
For every $u_{j}\in U$, we create a course $c_{j}$, with $O(c_{j})=1$. Let $C^{\prime}=\{c_{j}:1\leq j\leq n_{1}\}$. Additionally, we create two courses $e$ and $f$, with $O(e)=O(f)=2$. The upper quota of every course in $C^{\prime}\cup D^{2}$ is $1$, the upper quota of every course in $B\cup D^{1}\cup\{f\}$ is $2$, and the upper quota of $e$ is $n_{2}-K$. The lower quota of every course in $C^{\prime}\cup D^{2}\cup\{e\}$ is 0, whilst the lower quota of every course in $B\cup D^{1}\cup\{f\}$ is $2$.
We also create $K$ students $P=\{p_{k}:1\leq k\leq K\}$ such that, for every $p_{k}\in P$, $T(p_{k})=1$, and we create two students $q$ and $r$ such that $T(q)=T(r)=2$.
The master list of students ($\succ_{MLS}$) is
$s_{1}^{1}\succ_{MLS}s_{1}^{4}\succ_{MLS}s_{1}^{2}\succ_{MLS}s_{2}^{1}\succ_{MLS}s_{2}^{4}\succ_{MLS}s_{2}^{2}\succ_{MLS}...\succ_{MLS}s_{n_{2}}^{1}\succ_{MLS}\\ s_{n_{2}}^{4}\succ_{MLS}s_{n_{2}}^{2}\succ_{MLS}s_{1}^{3}\succ_{MLS}s_{2}^{3}\succ_{MLS}...\succ_{MLS}s_{n_{2}}^{3}\succ_{MLS}\\ p_{1}\succ_{MLS}...\succ_{MLS}p_{K}\succ_{MLS}q\succ_{MLS}r$ (3)
@@
That is, the first students belong to subsets $S^{1}$, $S^{4}$ and $S^{2}$, such that every $s_{i}^{1}$ is followed by $s_{i}^{4}$ and then by $s_{i}^{2}$. Then we place students from $S^{3}$, and after them students from $P$, in all cases in increasing subscript order. Finally, we place students $q$ and $r$. The master list of courses ($\succ_{MLC}$) is
$b_{1}^{1}\succ_{MLC}b_{1}^{2}\succ_{MLC}b_{2}^{1}\succ_{MLC}b_{2}^{2}\succ_{MLC}...\succ_{MLC}b_{n_{2}}^{1}\succ_{MLC}b_{n_{2}}^{2}\succ_{MLC}\\ d_{1}^{1}\succ_{MLC}d_{1}^{2}\succ_{MLC}d_{2}^{1}\succ_{MLC}d_{2}^{2}\succ_{MLC}...\succ_{MLC}d_{n_{2}}^{1}\succ_{MLC}d_{n_{2}}^{2}\succ_{MLC}\\ c_{1}\succ_{MLC}c_{2}\succ_{MLC}...\succ_{MLC}c_{n_{1}}\succ_{MLC}e\succ_{MLC}f$ (4)
@@
That is, the first courses are those from $B$, such that every $b_{i}^{1}$ is followed by $b_{i}^{2}$, then the next courses are from $D^{1}\cup D^{2}$, such that every $d_{i}^{1}$ is followed by $d_{i}^{2}$; after these, the courses in $C^{\prime}$ follow, in all cases in increasing subscript order. Finally we have $e$ followed by $f$.
@@
If in the exact-mm instance $w_{i}$ is matched with $u_{j}$, then $s_{i}^{1}$ is to be matched with $c_{j}$ and $b_{i}^{1}$, $s_{i}^{2}$ with $d_{i}^{1}$, $s_{i}^{3}$ with $d_{i}^{2}$, and $s_{i}^{4}$ with $d_{i}^{1}$. Similarly if $w_{i}$ is matched with $u_{k}$, then $s_{i}^{1}$ is to be matched with $d_{i}^{1}$, $s_{i}^{2}$ with $c_{k}$ and $b_{i}^{2}$, $s_{i}^{3}$ with $d_{i}^{2}$ and $s_{i}^{4}$ with $d_{i}^{1}$. If, however, $w_{i}$ is unmatched, $s_{i}^{1}$ is to be matched with $d_{i}^{1}$, $s_{i}^{2}$ with $d_{i}^{2}$, $s_{i}^{3}$ with $e$, and $s_{i}^{4}$ with $d_{i}^{1}$. Student $s_{i}^{3}$ contains courses $c_{j}$ and $c_{k}$ in her preference list to make sure that a course-complete pair-stable matching in $I$ corresponds to a maximal matching in $G$, so that if $s_{i}^{3}$ is matched with $e$, some students from $S^{1}$ or $S^{2}$ are matched with $c_{j}$ and $c_{k}$. The preference lists of students in $P$ includes every course in $B$.
@@
Students from $P$ are matched to the courses in $B$ that have some $s_{i}^{l}$ matched to them, which correspond to vertices $w_{i}\in G$ that are matched in $G$. The preference lists of $q$ and $r$ are:
@@
The preference list of every $c_{j}\in C^{\prime}$ depends on the vertices adjacent to $u_{j}$. For every $w_{i}$ adjacent to $u_{j}$, $c_{j}$ includes in its preference list either $s_{i}^{1}$ or $s_{i}^{2}$, as well as $s_{i}^{3}$. The order of all students is derived from the master list. If, for example, $u_{j}$ is adjacent to $w_{3}$, $w_{4}$ and $w_{5}$, the preference list of $c_{j}$ is given below. Here the superscript $l_{i}$ indicates a number that can be either $1$ or $2$. If $u_{j}$ is matched in $G$ then $c_{j}$ is matched with some student from $S^{1}$ or $S^{2}$, otherwise it is unmatched.
@@
Figure 8: Example of an exact-$m$ instance, with $K=2$, and some of the preference lists in the transformed instance of ca-ml-lq-cl.
We now show that a maximal matching $M$ of size $K$ in $G$ implies a student-complete pair-stable matching $M^{\prime}$ in $I$ that respects lower quotas. For every pair $(u_{j},w_{i})\in M$, we add to $M^{\prime}$ either the pairs $(s_{i}^{1},c_{j})$ and $(s_{i}^{1},b_{i}^{1})$, or the pairs $(s_{i}^{2},c_{j})$ and $(s_{i}^{2},b_{i}^{2})$, depending on whether $s_{i}^{1}$ or $s_{i}^{2}$ has $c_{j}$ in her preference list. If $(s_{i}^{1},c_{j})\in M^{\prime}$, then we add $(s_{i}^{2},d_{i}^{1})$ to $M^{\prime}$, otherwise we add $(s_{i}^{1},d_{i}^{1})$ to $M^{\prime}$. We also add $(s_{i}^{3},d_{i}^{2})$ and $(s_{i}^{4},d_{i}^{1})$ to $M^{\prime}$. If $w_{i}$ has no partner in $M$, we add the following pairs to $M^{\prime}$: $(s_{i}^{1},d_{i}^{1})$, $(s_{i}^{2},d_{i}^{2})$, $(s_{i}^{3},e)$, and $(s_{i}^{4},d_{i}^{1})$. Additionally, let $j_{k}$ ($1\leq k\leq K$) be a sequence such that $w_{j_{k}}$ is matched in $M$. For each $k$ ($1\leq k\leq K$), if $(s_{i}^{1},c_{j_{k}})\in M^{\prime}$ for some $s_{i}^{1}$, add $(p_{j_{k}},b_{i}^{1})$ to $M^{\prime}$, otherwise add $(p_{j_{k}},b_{i}^{2})$ to $M^{\prime}$. Finally, add pairs $(q,f)$ and $(r,f)$ to $M^{\prime}$. We close every course in $B$ that has no students matched to it. The matching $M^{\prime}$ is student-complete, because for every student $s\in S$, $T(s)=O(C_{M^{\prime}}(s))$. It also respects the lower quotas of every course that appears in $M^{\prime}$: in particular, course $f$ has two students, and every open course in $B$ and $D^{1}$ has two students also. We show below that $M^{\prime}$ is pair-stable.
@@
$M^{\prime}$ is a pair-stable matching in $I$.
@@
We prove the pair stability of $M^{\prime}$ by showing that no student can belong to a blocking pair. Every $s_{i}^{1}$ is matched with either $b_{i}^{1}$ (which is her first choice) and some $c_{j}$, or with $d_{i}^{1}$. In the first case, $s_{i}^{1}$ cannot block with $d_{i}^{1}$, as she cannot drop enough credits to do so. In the second case, $b_{i}^{1}$ is closed, so she cannot form a blocking pair with $b_{i}^{1}$. Similarly, every $s_{i}^{2}$ is matched with either (i) $b_{i}^{2}$ (her first choice) and some $c_{k}$, (ii) with $d_{i}^{1}$, or (iii) with $d_{i}^{2}$. If $s_{i}^{2}$ is matched with $b_{i}^{2}$ and $c_{k}$, or with $d_{i}^{1}$, the same reasoning applies as for $s_{i}^{1}$, and if she is matched with $d_{i}^{2}$, then $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$ and $(s_{i}^{4},d_{i}^{1})\in M^{\prime}$, so there is no blocking pair containing $s_{i}^{2}$. Student $s_{i}^{4}$ is always matched with the only course in her preference list, $d_{i}^{1}$.
@@
We now consider $s_{i}^{3}$. This student is matched with either $d_{i}^{2}$, her most-preferred course, or with $e$. In the second case, $(s_{i}^{2},d_{i}^{2})\in M^{\prime}$ and $w_{i}$ is unmatched in $M$. As $M$ is maximal in $G$, $u_{j}$ and $u_{k}$ are matched in $M$ with some $w_{p}$ and $w_{q}$ respectively. Thus, there exist $s_{p}^{l_{p}},s_{q}^{l_{q}}\in S^{1}\cup S^{2}$ such that $(s_{p}^{l_{p}},c_{j})\in M^{\prime}$ and $(s_{q}^{l_{q}},c_{k})\in M^{\prime}$. Each of these courses prefers its assigned student in $M^{\prime}$ to $s_{i}^{3}$. Hence, $s_{i}^{3}$ cannot form a blocking pair with any other course. Moving on to students in $P$, every $p_{k}\in P$ is matched with some $b_{i}^{r}\in B$, where $r\in\{1,2\}$. Let $b_{j}^{s}$ be an open course that $p_{k}$ prefers over $b_{i}^{r}$. Then $b_{j}^{s}$ is matched with both some $s_{l}^{1}$ or $s_{l}^{2}$, and also with a $p_{k}^{\prime}$ such that $p_{k}^{\prime}\succ_{MLS}p_{k}$. Thus, $p_{k}$ cannot form a blocking pair with $b_{i}^{l}$. Finally, $r$ is matched with the only course in her preference list, $f$, and $q$ is matched with $f$, too. Even though $q$ prefers $e$ over $f$, $e$ has filled its capacity, and every student enrolled in $e$ belongs to $S^{3}$, all of which appear in the master list before $q$. Therefore, no student belongs to a blocking pair. ∎
@@
Now, assume that there exists a student-complete pair-stable matching $M^{\prime}$ in $I$. We show how to construct a maximal matching $M$ of size $K$ in $G$. First, as $M^{\prime}$ is student-complete, $r$ is matched with $f$. The minimum quota of $f$ is $2$, so we have that $(q,f)\in M^{\prime}$ too. However, $q$ prefers $e$ over $f$. Thus, because $M^{\prime}$ is pair-stable, $e$ must have $n_{2}-K$ students assigned to it that are better than $q$. Hence, there are $n_{2}-K$ students from $S^{3}$ matched to $e$ in $M^{\prime}$. As $M^{\prime}$ is student-complete, $(s_{i}^{4},d_{i}^{1})\in M^{\prime}$, for all $i$ ($1\leq i\leq n_{2}$) and thus every course in $D^{1}$ is open. In order to preserve pair stability, for each $s_{i}^{3}$ matched to $e$, $M^{\prime}$ contains the pairs $(s_{i}^{1},d_{i}^{1})$ and $(s_{i}^{2},d_{i}^{2})$.